OpenStudy (anonymous):

ratios and proprtions: if you go to chicago and use 2/5 of a gallon tank of gas, how many gallons did you use? this was an 80-mile trip, how many tanks would you use to travel 500 miles?

OpenStudy (cwtan):

x = Full tank (2/5)x=80 mile x=200mile y= How many tank? xy=500 200y=500 y=2.5 tanks

OpenStudy (anonymous):

amount of gallons of gas used for chicago trip=2/5*1 gallon= 0.4 this was for 80 miles..For 1 miles 0.4/80 so for 500 miles =0.4/80*500